1.Consider the Solow model with exogenous growth. Assume that because of global warming the depreciation rate increases. Illustrate the change in the steady state. What happens to the growth rate of standard of living in the new steady state?
Answer #1

y=f(R) 82>S, loi Y2 (8z+n)k 6,n) k gizio Elke E i=xf(R) k2 ki k

Initial Steady state occurs at point E where the investment per worker curve (i = s*f(k)) intersect the straight depreciation line ((\delta1+n)*k).

The steady state per worker capital and output are k1 and y1.

Increase in the depreciation rate will shift the depreciation line upward to (\delta2+n)*k where \delta 2 > \delta 1.

Due to increase in depreciation there is a backward movement along the investment per worker curve from E to E1.

At this new steady state, both per worker capital and output reduced to k2 and y2.

Standard of living is measured by the stock of output per worker. Due to increase in depreciation, there is a fall in the output per worker from y1 to y2.

Growth rate of standard of living in the new steady is constant as the output growth rate = 0